OpenStudy (anonymous):

Simplify 2(x-3)^2 -7

OpenStudy (anonymous):

First, you want to solve (x-3)^2. Which is just (x-3)(x-3). You multiply the first term in the first parentheses with everything in the other, then the same with the second term. You should get: x^2-6x+9. Now going back to the original equation, you'd have 2+x^2-6x+9 = x^2-6x+2+9= x^2-6x+11

OpenStudy (anonymous):

Sorry, made a little mistake: Now going back to the original equation, you'd have 2+x^2-6x+9-7 = x^2-6x+2+9-7= x^2-6x+4

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

no and thankyou very much

OpenStudy (anonymous):

No problem.